OpenStudy (anonymous):

What is the simplified form of 3/4x+3 + 21/8x^2-14x-15 ?

OpenStudy (campbell_st):

well you'll need to factor the 2nd denominator \[\frac{3}{4x + 3} + \frac{21}{(2x - 5)(4x + 3)}\] you'll need a common denominator of \[(2x -5)(4x + 3)\] so multiply the 1st term by \[\frac{(2x -5)}{(2x -5)} \] and then add which will give \[\frac{3}{(4x + 3)} \times \frac{(2x -5)}{2x -5)} + \frac{21}{(2x -5)(4x + 3)}\] just simplify for your answer.

OpenStudy (anonymous):

6(2x+5)/(2x-5)(4x+3) ?

OpenStudy (campbell_st):

not quite you need to check the numerator I think its 6x - 15 + 21 = 6x + 6 = 6( x+1)